For the following chemical reaction, what mass of phosphoric acid, H3PO4, in grams will be needed to produce 1.71 mol of water, H2O?

3KOH+H3PO4⟶K3PO4+3H2O
1
Expert's answer
2019-02-15T02:22:21-0500

Chemical reaction:

3KOH + H3PO4 ⟶ K3PO4 + 3H2O

Mr (H3PO4)= 3(Ar(H)) + Ar(P) + 4(Ar(O))= 3 × 1+ 31 + 4 × 16 = 98 g/mol

Mr = M = 98 mol

m(H3PO4) = x = (98 mol × 1.71 mol) / (3 mol) = 55,86 g.

Answer: Mass of phosphoric acid, H3PO4 is 55,86 g.